A particular brand of coffee contains an average of 112 mg of caffeine per cup with a standard deviation of 20 mg. A barista wants to investigate the same to estimate the true population mean caffeine content correct to within 5 mg adopting 95% confidence. How many cups of the same brand of coffee does he need for a sample?


1
Expert's answer
2021-06-22T17:45:05-0400

"n=({\frac{\sigma z_{\alpha/2}}{SE}})^2"

"=(\frac{1.96*20}{5})^2"

=61.46

=61 cups